=== 2018/10/18(Thur.) ===

==== Attendance ====
 * Control room: Miyakawa, Yamamoto, Yuzurihara, Miyo, nakano, Tanaka, Yokozawa, Oshino, Dankan, Lucia, Fujii, Yokogawa, Enomoto, Kokeyama
 * Fan hut:
 * eZuce: 
 * Zoom: Hayakawa

==== Laser hazard ====

||||                        ||Before 3PM||After 3PM ||
||||                        ||          ||||
||Center||PSL               ||IR & Green||IR & Green||
||Center||around PR2        ||IR & Green||IR & Green||
||Center||behinde SR2&Yfront||IR        ||IR        ||
||||X end 1F                ||IR & Green||IR & Green||
||||Y end 1F                ||N/A       ||N/A       ||

==== Night workers ====

 * MIF: Enomoto, Yokogawa, Kozakai

=== Weekend ===


==== Transition from morning work (*:responsible person) ====

 * IPs of ETMX are freely moving. Oplev at ETMX has still no light.
 * ITMX alignment seems to be fine already. We can see green light at REFL port.
 * Leak test is going on.

==== Report of work (*:responsible person) ====
 * KAGRA summary pages are available now.

==== Work plan (*:responsible person) ====

 * Green is hitting ETM. Alignment of ITM will be done.
 * Needs damping of ITMX and ETMX.
 * Marionet oplev light are not existing for both ITMX/ETMX.